OpenStudy (anonymous):

anybody know what fraction equals 27/8 and 1 1/8 (btw 1 1/8 is a mixed number)

OpenStudy (anonymous):

you want to rewrite 27/8 as a mixed number?

OpenStudy (anonymous):

i ment 1 and 1/8 is a mixed number

OpenStudy (anonymous):

1 1/8 + 27/8= Change the mixed number to an improper fraction. So 1 1/8 becomes 9/8 Now just add 9/8+27/8= 36/8 Reduce 36/8 to 4 1/2